OpenStudy (anonymous):

How many moles of Pb(NO3)2 are there in 10.0 mL of a 0.10 M solution?

Parth (parthkohli):

Hi! 0.10 M means that the number of moles of Pb(NO3)2 divided by the number of liters of the solution is 0.10.

Parth (parthkohli):

\[0.10=\dfrac{\rm no~ of ~moles ~of~Pb(NO_3)_2 }{1/100}\]

Parth (parthkohli):

Note that the 1/100 is there because 10 mL = 1/100 of a litre. Now solve the equation.

OpenStudy (anonymous):

to find the number of moles of Pb(NO3)2, I first find the molar mass which comes to 331, and then how do I convert from g and ml to moles?

Parth (parthkohli):

No... you don't need to do that. I gave an equation above. Take a look at it.

OpenStudy (anonymous):

I don't understand how to solve without a value for the numerator

Parth (parthkohli):

You want to get the number of moles of Pb(NO3)2. The equation is\[0.10=\dfrac{\rm no~ of ~moles ~of~Pb(NO_3)_2 }{1/100}\]

Parth (parthkohli):

Take 1/100 to the other side.